OpenStudy (anonymous):

How to prove this identity: 1-sin2x/cosx = cosx?

OpenStudy (anonymous):

do I expand sin2x? eep.

OpenStudy (lgbasallote):

is that sin (2x) or sin^2 x?

OpenStudy (anonymous):

sin^2x

OpenStudy (lgbasallote):

lemme rephrase that.. \[\frac{1 - \sin^2x}{\cos^2x}\]

OpenStudy (anonymous):

yup

OpenStudy (lgbasallote):

i mean cos x

OpenStudy (anonymous):

okay

OpenStudy (lgbasallote):

change \(1 - \sin^2x\) into one of the good old trig identities

OpenStudy (aravindg):

is that sin2x or sin^2x

OpenStudy (anonymous):

cos^2x?

OpenStudy (lgbasallote):

got it! \[\frac{\cos^2 x}{\cos x}\] simplify that

OpenStudy (lgbasallote):

lol mimi :P

OpenStudy (mimi_x3):

\(sin^2x+cos^2x=1\) => \(cos^2x=1-sin^2x\)

OpenStudy (anonymous):

ohhh, i cancel: cosxcosx/cosx?

OpenStudy (lgbasallote):

\[\frac{\cancel{(\cos x)}(\cos x)}{\cancel{\cos x}}\]